Think of something that you might want to measure that is affected by random variation. Identify what you want to measure, then describe its (approximate) sample space. Give a rough description of the probabilities associated with those values (you can simply specify if they are all the same probability or if values in one range will be more likely than values in another range). What would you say to a person who says that he or she "knows" what the outcome of an individual observation will be (an outcome of something that has not happened yet that is subject to random error)?

1
Expert's answer
2021-12-10T01:24:06-0500

For example, if we throw a coin then probabilities of events head, the tail will be "\frac{1}{2}, \frac{1}{2}" or if we roll a dice then the probability of getting any of 1, 2, 3, 4, 5 and 6 will be "\frac{1}{6}". Hence, in this example, both cases probability is equal for each event.

But if we have 4 white balls and 5 black balls in the container then the probability of drawing a white ball is "\frac{4}{9}" and probability of blackball is "\frac{5}{9}". Therefore in this example probability are not equal for a white and black ball.

Hence probability is affected by a random variation.